OpenStudy (anonymous):

(+24) – (-6) + (-8) – (-2) =

OpenStudy (anonymous):

8

OpenStudy (anonymous):

(+24) – (-6) + (-8) – (-2) = +24 – (-6) + (-8) – (-2) = 24 – (-6) + (-8) – (-2) = 24 + 6 + (-8) – (-2) = 24 + 6 -8 – (-2) = can you do the rest now?

OpenStudy (anonymous):

(+24) – (-6) + (-8) – (-2) =24

OpenStudy (anonymous):

correct ^_^

OpenStudy (anonymous):

lol I didnt see the add mark there